OpenStudy (anonymous):

7. Solve the exponential equation. 125^9x – 2 = 150 (1 point) –1.8847 –0.1069 0.3375 1.0378 8. Solve log(3x + 2) = 3. (1 point) 1000/3 998/3 998 1/3

OpenStudy (anonymous):

125^9x=152 then take log from both sides

OpenStudy (anonymous):

but what is log? how do I do that???

OpenStudy (anonymous):

125^9x=(5^3)^9x (27x)} \]